# Quantum Key Distribution via Charge Teleportation

### v1

@@ I Introduction
Focusing on this primary application, we develop and analyze a QKD protocol based on the teleportation of a local charge observable. We investigate the protocol’s performance under two distinct interaction models for the underlying many-body system: a star-coupled Transverse Field Ising Model (TFIM), $H^{(1)}=J\sum_{k=1}^{N}X_{0}X_{k}+h\sum_{k=0}^{N}Z_{k}$, where a central qubit (Alice) interacts with all others, and a 1D nearest-neighbor TFIM, $H^{(2)}=J\sum_{k=1}^{N}X_{k-1}X_{k}+h\sum_{k=0}^{N}Z_{k}$. Through numerical simulations and Qiskit-based quantum circuit analysis, we compare the efficacy and robustness of charge teleportation against energy teleportation for these models. Our findings show that charge teleportation exhibits superior symmetry, stability against statistical noise, and scalability with system size, making it a more promising candidate for practical QKD implementations. Furthermore, we analyze the protocol’s resilience to various noise channels, identifying the operational regimes where secure key distribution remains viable.
@@ II.1.1 Protocol Steps
    Bob’s Conditional Rotation: Based on the classical information received, Bob applies a conditional rotation on his subsystem at site $N$:
@@ II.2 QKD with Charge teleportation
We now apply the general observable teleportation protocol to the charge operator. The charge at Bob’s site (site $N$) is defined as:
@@ III.1.1 Star-Coupled TFIM Hamiltonian Definition
where the global state resides in the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes(N+1)}$. The site 0 is assigned to Alice, while site $N$ is assigned to Bob. The corresponding localized Hamiltonian on Bob’s side, derived from the global Hamiltonian by isolating terms involving site $N$, is:
@@ III.2.1 Nearest Neighbors Hamiltonian
where, again, site 0 is assigned to Alice and site $N$ to Bob. The Hilbert space is $\mathcal{H}=(\mathbb{C}^{2})^{\otimes(N+1)}$. The localized Hamiltonian for Bob’s site (site $N$) includes the interaction with its nearest neighbor and is given by:
@@
For a system with $2$ sites, the global Hamiltonian $H^{(2)}$ reduces to:
@@ III.3.1 Feasibility of Analytical Solutions
As discussed earlier, the Hilbert space for the system scales as $2^{N+1}$, making analytical treatment rapidly intractable as $N$ increases. Specifically, for $N=2$ (i.e., a 3-site system), one must diagonalize an $8\times 8$ Hamiltonian matrix. Although symmetry considerations allow decomposition into smaller blocks—e.g., a $2\times 2$ and a $6\times 6$ block—there is no closed-form solution for the ground state for arbitrary values of $h$ and $J$ when $N\geq 2$.
@@ IV.1.1 Teleportation Behavior vs Coupling Strength $J$
  • Scalability of extracted signal— Importantly, since both energy and charge observables were rescaled to arbitrary units, we may compare their relative magnitudes. Our simulations reveal that the charge expectation value at Bob’s site remains approximately constant as the system size $N$ increases, whereas the extracted energy diminishes. This suggests that charge teleportation may be more favorable for large-scale implementations, especially in distributed QKD architectures. This suggests that charge teleportation may be more favorable for large-scale implementations, especially in distributed QKD architectures. Furthermore, the perfect symmetry of the charge signal is a decisive advantage for cryptography; it ensures that the logical ’0’ and ’1’ are equally distinguishable, preventing any potential bias in the raw key that an eavesdropper might exploit.
@@ IV.2.1 Protocol Implementation and Basis Choices
Interestingly, while both energy and charge teleportation degrade as $N$ increases, the energy protocol appears more robust to the choice of measurement basis. This is clearly seen in Figure 4, where the curves corresponding to different bases remain relatively close in the energy case but diverge significantly in the charge case.
@@ V.2 Star-Coupled Hamiltonian Results
As system size $N$ increases, the statistical noise in the energy teleportation results becomes more pronounced. This effect is especially visible in panel (c), where the observable fluctuations for $N=3$ exhibit substantial variance due to the low signal-to-noise ratio of energy teleportation at larger $N$. Charge teleportation, in contrast, remains smooth and symmetric across all system sizes (Fig. 6, bottom row), highlighting its superior robustness. This stands in clear opposition to the energy results, which show significant statistical fluctuations (Fig. 6, top row).
@@ Discussion
  • Propagation through intermediate sites: In the nearest-neighbors model, Alice and Bob are no longer directly coupled. The teleportation signal must propagate through a chain of intermediate qubits, weakening the effect as $N$ increases.
@@ V.5 Summary: Insights from Qiskit Simulations
  • Charge vs. Energy Teleportation: Across all experiments, charge teleportation demonstrates higher stability, cleaner signal separation, and reduced statistical noise. This superior statistical clarity, stemming from its single-operator measurement, is a key practical advantage for QKD. It implies that a charge-based protocol would require fewer experimental shots to establish a key bit with high confidence, making it more efficient and scalable with system size $N$ on near-term quantum devices.
  • Hardware Scalability: While teleportation works well for small $N$, scaling to larger systems introduces challenges. In the nearest-neighbor model, indirect interaction between Alice and Bob weakens the effective signal and increases the number of shots required for convergence. This highlights the trade-off between physical model realism and practical observability.
@@ VI.1 Classical Communication Error
The numerical results in Figure 9 show this linear decay for all simulated configurations. For the star-interaction model $H^{(1)}$, a clear sign-flip threshold appears around $p\approx 0.25$, consistent with the findings in [QKDbyQET]. While the absolute magnitude of the energy signal is larger than that of the charge signal in this model, both are similarly vulnerable to this error (see Appendix Fig. 24 and Fig. 25 for a direct comparison). However, for the more realistic nearest-neighbor model ($H^{(2)}$), the energy signal diminishes rapidly with system size $N$, whereas the charge signal remains more stable (a trend clearly visible in the energy data in Appendix Fig. 23). This highlights that for larger distances, the charge protocol’s signal is more resilient, even though the fundamental error threshold remains similar.
@@ VI.4 Bit-Flip and Phase-Flip Errors
This error, modeled by $\rho\rightarrow(1-p)\rho+pX_{n}\rho X_{n}$, degrades the specific quantum correlations leveraged by the protocol. As shown in Figure 13, both energy and charge signals decay smoothly with the error probability $p$. Sign-crossings are rare, indicating that the protocol is relatively robust against this error type. Energy generally maintains a larger absolute signal, which in principle makes it easier to measure, provided the noise level is low enough to prevent ambiguity (cf. the energy data in Appendix Figs. 3739, and 40).
@@
Note. Additional configurations, including extended $N$ sweeps and further comparisons, are introduced in Appendix F for completeness.
@@ VII.3 Parameter Estimation
Figure 15: Asymptotic key rate analysis for the $H^{(2)}$ model ($J=2,h=1$) under phase-flip noise on Bob’s site. Plots show the $Q_{BER}$ ($e_{\text{bit}}$), the test basis error ($e_{\text{ph}}$), and the resulting asymptotic key rate ($K_{\text{asym}}$) as a function of noise probability $p$.
@@
  • Test Basis ($e_{\text{ph}}$): We run a separate simulation in the test basis ($\sigma_{A}=Y_{0}$) with the same noise $p$. We calculate the analogous probabilities $P^{\prime}_{+}(p)$, $P^{\prime}_{-}(p)$, and $P^{\prime}_{0}(p)$ for the $\Delta Q_{B}$ measurement. As with the key basis, the $P^{\prime}_{0}(p)$ events are sifted out. The test-basis error $e_{\text{test}}(p)=P^{\prime}_{+}(p)/(P^{\prime}_{+}(p)+P^{\prime}_{-}(p))$ (orange line) provides the upper bound for the phase error rate, so we set $e_{\text{ph}}(p)\approx e_{\text{test}}(p)$.
@@
The results for $N=2,3,$ and $4$ (with $J=2,h=1$) are presented in Figure 15.
The simulation results in Fig. 15 reveal a critical dependence on the system size $N$. For the minimal chain $N=2$ (Fig. 15a), the protocol is secure. The asymptotic key rate $K_{\text{asym}}$ (green line) is positive for zero noise, $K_{\text{asym}}(p=0)>0$, indicating a secure channel. This is achieved despite a high intrinsic phase error $e_{\text{ph}}(p=0)\approx 0.19$ because the bit error rate is sufficiently low, $e_{\text{bit}}(p=0)\approx 0.04$. The resulting key rate $K_{\text{asym}}(p=0)=1-h(0.04)-h(0.19)\approx 0.05$ is small but positive. The protocol can tolerate noise up to a threshold of $p_{\text{th}}\approx 0.02$, after which $K_{\text{asym}}$ drops to zero.
@@ VII.5 Finite-Key Effects and Future Work
Our analysis confirms the fundamental viability of the charge teleportation mechanism for QKD. The next steps involve a) optimizing the Hamiltonian parameters $J$ and $h$ to maximize the intrinsic key rate and secure distance $N$, and b) applying finite-key analysis to the optimized protocol to determine the practical key rates achievable with a realistic number of signals.
@@ VIII Summary
Our primary finding is the superior stability and robustness of the charge-based protocol. While energy teleportation can, in certain regimes, yield a larger absolute signal ($\langle\Delta H_{B}\rangle$), it is highly sensitive to statistical noise and exhibits an asymmetric response to Alice’s classical bit choice. In contrast, the teleported charge expectation value ($\langle\Delta Q_{B}\rangle$) shows a perfect mirror symmetry, which is crucial for a reliable bit-to-key mapping. This symmetry, combined with a signal that remains stable and scalable as the system size $N$ increases, makes the charge protocol inherently more resilient to the finite-shot noise that characterizes near-term quantum hardware.
@@ B.1.1 System Definition and Local Hamiltonian
The localized Hamiltonian at Bob’s site (site $N$) is obtained by isolating the terms acting on site $N$:
@@ C.2.2 Summary of Observations
  • The charge protocol remains strongly dependent on the basis choice, with the difference between X and Y observables increasing with $N$.
@@ E.4 Summary and Conclusions
  • The ”meaningful statistical noise” observed for larger systems (e.g., $N=3,4$ in the nearest-neighbor model) is a result of a decreasing Signal-to-Noise Ratio (SNR). As $N$ increases, the teleported signal $|\langle O\rangle|$ diminishes due to weaker correlations, while the single-shot variance $\sigma_{O}^{2}$ remains approximately constant (approaching 1). Consequently, the SEM $\sigma_{\langle O\rangle}$ becomes comparable to or larger than the signal itself, requiring a significantly larger $n_{\text{shots}}$ to resolve.
@@ F.1 Classical Communication Error
Figure 23: Energy vs. Classical communication error for different NN values. The linear degradation of the signal is evident across all models, with a consistent failure threshold where the expectation value crosses zero. The signal magnitude for the $H^{(2)}$ model visibly diminishes with increasing system size $N$.
@@
Figure 23 shows the teleported energy as a function of the error probability $p$ for various system sizes $N$. For all Hamiltonians, the expected linear decay is observed. A critical failure threshold, where the sign of the teleported energy flips, consistently appears around $p\approx 0.25$, corroborating the analysis in [QKDbyQET]. Notably, for the nearest-neighbor model ($H^{(2)}$), the magnitude of the teleported energy signal weakens significantly as the system size $N$ increases, which makes resolving the signal from noise more challenging in larger systems.
@@
  • Nearest-Neighbor Model ($H^{(2)}$): The results for this model, depicted in Figures 25 ($\sigma_{A}=X_{0}$) and 26 ($\sigma_{A}=Y_{0}$), confirm a similar linear degradation. However, a key distinction emerges: the charge signal’s magnitude remains more robust against increases in system size $N$ compared to the energy signal. While the fundamental error threshold remains in the same $p\approx 0.25-0.32$ range, the rapid decay of the energy signal’s absolute value means that for larger distances between Alice and Bob, the charge-based protocol provides a more resilient and discernible signal. This highlights a significant practical advantage of charge teleportation for implementing QKD in larger, more realistic quantum systems.
@@ F.3 Superposition with an Excited State
where $p$ is the probability of the excited state amplitude and $\alpha$ is a relative phase. As shown in [QKDbyQET], the effect of the phase $\alpha$ is generally negligible, so the analysis focuses on the impact of the probability $p$.
@@
The trend, shown across Figures 32 to 36, is qualitatively similar to the statistical mixture case but quantitatively distinct due to the non-linear signal degradation. Once again, the energy protocol proves highly susceptible to this error, with the signal quickly flipping to positive at low error probabilities. The charge protocol, however, demonstrates significant resilience. The correlations enabling charge teleportation are less affected by this coherent mixing, allowing the signal to maintain its negative sign and thus preserve the integrity of the key bit across a much wider range of $p$.

### v2

@@ I Introduction
Focusing on this primary application, we develop and analyze a QKD protocol based on the teleportation of a local charge observable. We investigate the protocol’s performance under two distinct interaction models for the underlying many-body system: a star-coupled Transverse Field Ising Model (TFIM), $H^{(1)}=J\sum_{k=1}^{N}X_{0}X_{k}+h\sum_{k=0}^{N}Z_{k}$, where a central qubit (Alice) interacts with all others, and a 1D nearest-neighbor TFIM, $H^{(2)}=J\sum_{k=1}^{N}X_{k-1}X_{k}+h\sum_{k=0}^{N}Z_{k}$. Through numerical simulations and Qiskit-based quantum circuit analysis [Amir_QKD_by_Charge_Teleportation], we compare the efficacy and robustness of charge teleportation against energy teleportation for these models. Our findings show that charge teleportation exhibits superior symmetry, stability against statistical noise, and scalability with system size, making it a more promising candidate for practical QKD implementations. Furthermore, we analyze the protocol’s resilience to various noise channels, identifying the operational regimes where secure key distribution remains viable.
@@ II.1.1 Protocol Steps
    Bob’s Conditional Rotation: Based on the classical information received, Bob applies a conditional rotation on his subsystem at site NN:
@@ II.2 QKD with Charge teleportation
We now apply the general observable teleportation protocol to the charge operator. The charge at Bob’s site (site NN) is defined as:
@@ III.1.1 Star-Coupled TFIM Hamiltonian Definition
where the global state resides in the Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes(N+1)}$. The site 0 is assigned to Alice, while site NN is assigned to Bob. The corresponding localized Hamiltonian on Bob’s side, derived from the global Hamiltonian by isolating terms involving site NN, is:
@@ III.2.1 Nearest Neighbors Hamiltonian
where, again, site 0 is assigned to Alice and site NN to Bob. The Hilbert space is $\mathcal{H}=(\mathbb{C}^{2})^{\otimes(N+1)}$. The localized Hamiltonian for Bob’s site (site NN) includes the interaction with its nearest neighbor and is given by:
@@
For a system with 22 sites, the global Hamiltonian $H^{(2)}$ reduces to:
@@ III.3.1 Feasibility of Analytical Solutions
As discussed earlier, the Hilbert space for the system scales as $2^{N+1}$, making analytical treatment rapidly intractable as NN increases. Specifically, for $N=2$ (i.e., a 3-site system), one must diagonalize an $8\times 8$ Hamiltonian matrix. Although symmetry considerations allow decomposition into smaller blocks—e.g., a $2\times 2$ and a $6\times 6$ block—there is no closed-form solution for the ground state for arbitrary values of $h$ and $J$ when $N\geq 2$.
@@ IV.1.1 Teleportation Behavior vs Coupling Strength $J$
  • Scalability of extracted signal— Importantly, since both energy and charge observables were rescaled to arbitrary units, we may compare their relative magnitudes. Our simulations reveal that the charge expectation value at Bob’s site remains approximately constant as the system size NN increases, whereas the extracted energy diminishes. This suggests that charge teleportation may be more favorable for large-scale implementations, especially in distributed QKD architectures. This suggests that charge teleportation may be more favorable for large-scale implementations, especially in distributed QKD architectures. Furthermore, the perfect symmetry of the charge signal is a decisive advantage for cryptography; it ensures that the logical ’0’ and ’1’ are equally distinguishable, preventing any potential bias in the raw key that an eavesdropper might exploit.
@@ IV.2.1 Protocol Implementation and Basis Choices
Interestingly, while both energy and charge teleportation degrade as NN increases, the energy protocol appears more robust to the choice of measurement basis. This is clearly seen in Figure 4, where the curves corresponding to different bases remain relatively close in the energy case but diverge significantly in the charge case.
@@ V.2 Star-Coupled Hamiltonian Results
As system size NN increases, the statistical noise in the energy teleportation results becomes more pronounced. This effect is especially visible in panel (c), where the observable fluctuations for $N=3$ exhibit substantial variance due to the low signal-to-noise ratio of energy teleportation at larger NN. Charge teleportation, in contrast, remains smooth and symmetric across all system sizes (Fig. 6, bottom row), highlighting its superior robustness. This stands in clear opposition to the energy results, which show significant statistical fluctuations (Fig. 6, top row).
@@ Discussion
  • Propagation through intermediate sites: In the nearest-neighbors model, Alice and Bob are no longer directly coupled. The teleportation signal must propagate through a chain of intermediate qubits, weakening the effect as NN increases.
@@ V.5 Summary: Insights from Qiskit Simulations
  • Charge vs. Energy Teleportation: Across all experiments, charge teleportation demonstrates higher stability, cleaner signal separation, and reduced statistical noise. This superior statistical clarity, stemming from its single-operator measurement, is a key practical advantage for QKD. It implies that a charge-based protocol would require fewer experimental shots to establish a key bit with high confidence, making it more efficient and scalable with system size NN on near-term quantum devices.
  • Hardware Scalability: While teleportation works well for small NN, scaling to larger systems introduces challenges. In the nearest-neighbor model, indirect interaction between Alice and Bob weakens the effective signal and increases the number of shots required for convergence. This highlights the trade-off between physical model realism and practical observability.
@@ VI.1 Classical Communication Error
The numerical results in Figure 9 show this linear decay for all simulated configurations. For the star-interaction model $H^{(1)}$, a clear sign-flip threshold appears around $p\approx 0.25$, consistent with the findings in [QKDbyQET]. While the absolute magnitude of the energy signal is larger than that of the charge signal in this model, both are similarly vulnerable to this error (see Appendix Fig. 24 and Fig. 25 for a direct comparison). However, for the more realistic nearest-neighbor model ($H^{(2)}$), the energy signal diminishes rapidly with system size NN, whereas the charge signal remains more stable (a trend clearly visible in the energy data in Appendix Fig. 23). This highlights that for larger distances, the charge protocol’s signal is more resilient, even though the fundamental error threshold remains similar.
@@ VI.4 Bit-Flip and Phase-Flip Errors
This error, modeled by $\rho\rightarrow(1-p)\rho+pX_{n}\rho X_{n}$, degrades the specific quantum correlations leveraged by the protocol. As shown in Figure 13, both energy and charge signals decay smoothly with the error probability pp. Sign-crossings are rare, indicating that the protocol is relatively robust against this error type. Energy generally maintains a larger absolute signal, which in principle makes it easier to measure, provided the noise level is low enough to prevent ambiguity (cf. the energy data in Appendix Figs. 3739, and 40).
@@
Note. Additional configurations, including extended NN sweeps and further comparisons, are introduced in Appendix F for completeness.
@@ VII.3 Parameter Estimation
Figure 15: Asymptotic key rate analysis for the $H^{(2)}$ model ($J=2,h=1$) under phase-flip noise on Bob’s site. Plots show the $Q_{BER}$ ($e_{\text{bit}}$), the test basis error ($e_{\text{ph}}$), and the resulting asymptotic key rate ($K_{\text{asym}}$) as a function of noise probability pp.
@@
  • Test Basis ($e_{\text{ph}}$): We run a separate simulation in the test basis ($\sigma_{A}=Y_{0}$) with the same noise pp. We calculate the analogous probabilities $P^{\prime}_{+}(p)$, $P^{\prime}_{-}(p)$, and $P^{\prime}_{0}(p)$ for the $\Delta Q_{B}$ measurement. As with the key basis, the $P^{\prime}_{0}(p)$ events are sifted out. The test-basis error $e_{\text{test}}(p)=P^{\prime}_{+}(p)/(P^{\prime}_{+}(p)+P^{\prime}_{-}(p))$ (orange line) provides the upper bound for the phase error rate, so we set $e_{\text{ph}}(p)\approx e_{\text{test}}(p)$.
@@
The results for $N=2,3,$ and 44 (with $J=2,h=1$) are presented in Figure 15.
The simulation results in Fig. 15 reveal a critical dependence on the system size NN. For the minimal chain $N=2$ (Fig. 15a), the protocol is secure. The asymptotic key rate $K_{\text{asym}}$ (green line) is positive for zero noise, $K_{\text{asym}}(p=0)>0$, indicating a secure channel. This is achieved despite a high intrinsic phase error $e_{\text{ph}}(p=0)\approx 0.19$ because the bit error rate is sufficiently low, $e_{\text{bit}}(p=0)\approx 0.04$. The resulting key rate $K_{\text{asym}}(p=0)=1-h(0.04)-h(0.19)\approx 0.05$ is small but positive. The protocol can tolerate noise up to a threshold of $p_{\text{th}}\approx 0.02$, after which $K_{\text{asym}}$ drops to zero.
@@ VII.5 Finite-Key Effects and Future Work
Our analysis confirms the fundamental viability of the charge teleportation mechanism for QKD. The next steps involve a) optimizing the Hamiltonian parameters $J$ and $h$ to maximize the intrinsic key rate and secure distance NN, and b) applying finite-key analysis to the optimized protocol to determine the practical key rates achievable with a realistic number of signals.
@@ VIII Summary
Our primary finding is the superior stability and robustness of the charge-based protocol. While energy teleportation can, in certain regimes, yield a larger absolute signal ($\langle\Delta H_{B}\rangle$), it is highly sensitive to statistical noise and exhibits an asymmetric response to Alice’s classical bit choice. In contrast, the teleported charge expectation value ($\langle\Delta Q_{B}\rangle$) shows a perfect mirror symmetry, which is crucial for a reliable bit-to-key mapping. This symmetry, combined with a signal that remains stable and scalable as the system size NN increases, makes the charge protocol inherently more resilient to the finite-shot noise that characterizes near-term quantum hardware.
@@ B.1.1 System Definition and Local Hamiltonian
The localized Hamiltonian at Bob’s site (site NN) is obtained by isolating the terms acting on site NN:
@@ C.2.2 Summary of Observations
  • The charge protocol remains strongly dependent on the basis choice, with the difference between X and Y observables increasing with NN.
@@ E.4 Summary and Conclusions
  • The ”meaningful statistical noise” observed for larger systems (e.g., $N=3,4$ in the nearest-neighbor model) is a result of a decreasing Signal-to-Noise Ratio (SNR). As NN increases, the teleported signal $|\langle O\rangle|$ diminishes due to weaker correlations, while the single-shot variance $\sigma_{O}^{2}$ remains approximately constant (approaching 1). Consequently, the SEM $\sigma_{\langle O\rangle}$ becomes comparable to or larger than the signal itself, requiring a significantly larger $n_{\text{shots}}$ to resolve.
@@ F.1 Classical Communication Error
Figure 23: Energy vs. Classical communication error for different NN values. The linear degradation of the signal is evident across all models, with a consistent failure threshold where the expectation value crosses zero. The signal magnitude for the $H^{(2)}$ model visibly diminishes with increasing system size NN.
@@
Figure 23 shows the teleported energy as a function of the error probability $p$ for various system sizes NN. For all Hamiltonians, the expected linear decay is observed. A critical failure threshold, where the sign of the teleported energy flips, consistently appears around $p\approx 0.25$, corroborating the analysis in [QKDbyQET]. Notably, for the nearest-neighbor model ($H^{(2)}$), the magnitude of the teleported energy signal weakens significantly as the system size NN increases, which makes resolving the signal from noise more challenging in larger systems.
@@
  • Nearest-Neighbor Model ($H^{(2)}$): The results for this model, depicted in Figures 25 ($\sigma_{A}=X_{0}$) and 26 ($\sigma_{A}=Y_{0}$), confirm a similar linear degradation. However, a key distinction emerges: the charge signal’s magnitude remains more robust against increases in system size NN compared to the energy signal. While the fundamental error threshold remains in the same $p\approx 0.25-0.32$ range, the rapid decay of the energy signal’s absolute value means that for larger distances between Alice and Bob, the charge-based protocol provides a more resilient and discernible signal. This highlights a significant practical advantage of charge teleportation for implementing QKD in larger, more realistic quantum systems.
@@ F.3 Superposition with an Excited State
where $p$ is the probability of the excited state amplitude and $\alpha$ is a relative phase. As shown in [QKDbyQET], the effect of the phase $\alpha$ is generally negligible, so the analysis focuses on the impact of the probability pp.
@@
The trend, shown across Figures 32 to 36, is qualitatively similar to the statistical mixture case but quantitatively distinct due to the non-linear signal degradation. Once again, the energy protocol proves highly susceptible to this error, with the signal quickly flipping to positive at low error probabilities. The charge protocol, however, demonstrates significant resilience. The correlations enabling charge teleportation are less affected by this coherent mixing, allowing the signal to maintain its negative sign and thus preserve the integrity of the key bit across a much wider range of pp.